Solve (y-z/yz)p+(z-x/xy)q=x-y/yx


1
Expert's answer
2020-07-09T17:14:30-0400

The given problem must be, "\dfrac{y-z}{yz}p+\dfrac{z-x}{xz}q=\dfrac{x-y}{yx}" .


This problem can be written as, "x(y-z)p+y(z-x)q=z(x-y)".

The subsidiary equations are,

"\dfrac{dx}{x(y-z)} = \dfrac{dy}{y(z-x)} =\dfrac{dz}{z(x-y)}"


Using the multipliers 1,1,1 each of the above ratio is equal to

"\dfrac{dx+dy+dz}{x(y-z)+y(z-x)+z(x-y)}=\dfrac{dx+dy+dz}{0}"


Hence,

"\begin{aligned}
dx+dy+dz&=0,\\ \text{Integrating we get,}\\
x+y+z&=c_{1}
\end{aligned}"

Using multipliers "\frac{1}{x},\frac{1}{y},\frac{1}{z}" the ratio is equal to


"\dfrac{\frac{1}{x}dx+\frac{1}{y}dy+\frac{1}{z}dz}{(y-z)+(z-x)+(x-y)}=\dfrac{\frac{1}{x}dx+\frac{1}{y}dy+\frac{1}{z}dz}{0}"


Hence,

"\begin{aligned}
\frac{1}{x}dx+\frac{1}{y}dy+\frac{1}{z}dz&=0\\ \text{Integrating we get,}\\
\ln x+\ln y+\ln z&=\ln c_{2}\\
\ln (xyz)&=\ln c_{2}\\
xyz&=c_{2}
\end{aligned}"

Thus, the general solution is

"\begin{aligned}
\phi(c_{1},c_{2})&=0\\
\phi(x+y+z,xyz)&=0
\end{aligned}"